ABC Inc. recently issued $1,000 par bonds at a 5.25% coupon rate. The bonds have 15 years to maturity and current price of the bond is $850. If the call price is $1,050 and the bond can be called in 10 years, what is the yield to call? Assume semi-annual compounding.

Note: Convert your answer to percentage and round off to two decimal points. Do not enter % in the answer box.
Answer #1

Since the bond is callable in 10 years and we are required to calculate the YTC ( Yield to call), we take duration as 10 years and ignore the original maturity period of 15 years in the question.

Thus, we have ............

nper = number of periods = 10 * 2 semi annuals = 20

pmt = Coupon payment = 1000 * 0.0525 * 1/2 = 26.25

pv = Present value = current selling price = - 850 ( negative is needed as for buyer it cash out flow)

fv = Future value = 1050

Now, for computing YTC ............ use " RATE " function in excel

= Rate(nper, pmt, pv, fv )

= 3.8876% ............. But this represents semi annual rate. We have to express YTC in annual terms.

Final answer

YTC = 3.8876% * 2 = 7.78